# Performant Automatic BLAS Offloading on Unified Memory Architecture with OpenMP First-Touch Style Data Movement

### v1

@@ 4.3. Application Test 2: PARSEC
PARSEC (Pseudopotential Algorithm for Real-Space Electronic Calculations)(Kronik et al., 2006; Chelikowsky et al., 1994) is a package designed to perform Density Functional Theory (DFT) calculations of solids and molecules. It solves the Kohn–Sham equations directly in real space, avoiding the use of explicit basis sets. Our benchmark case calculates energy of a Silicon nanocrystal $Si_{1947}H_{604}$, boundary sphere radius is set to 50 bohr, grid spacing is 0.9 bohr, the calculation is limited to two self-consistent field steps to reduce benchmark cost, but performance characteristics of a fully converged calculation is identical.
@@
In these tests, the Mem-Copy data policy resulted in runtimes slower than dual-CPU execution. The cudaMemcpy operations consumed 220 seconds, accounting for more than 50% of the total runtime — a significantly higher proportion than observed in the MuST test case. This is due to the fact that most matrices used in PARSEC are long skinny matrices rather than squared shape ones. For example, a common dgemm input in PARSEC is transA=’T’, transB=’N’, M=32, N=2400, K=93536, the extreme skinny shapes make the total byte size of the matrices much bigger than if square matrices are used in a calculation with equivalent computational workload. a calculation of identical compute workload but with all square matrices. This outcome again demonstrates that the conventional data movement strategy for automatic offload is impractically useful even on Grace-Hoper where NVLink-C2C transfer rate is 450 GB/s per direction, not to mention the PCIe based cards where PCIe Gen5 x16 can only do 64 GB/s. The counter-based data migration strategy performs even worse due to incompetent migration algorithm, see complete discussions of NVIDIA’s migration issues later in Section 4.4.1.
Finally, the Device First-Use policy is able to efficiently manage the data movement, enabling significant performance gains. The total runtime is nearly 2x faster than the CPU run, with dgemm component achieving nearly 10x speedup compared to CPU run. Data transfer overhead is minimal, totaling just 1.3 seconds. The speedup is able to offset the extra cost from power hungry GPU. Again, data reuse is counted and for every matrix that is migrated to GPU resident memory, it is reused on average 570 times by subsequent dgemm calls.
@@ 4.4. Performance Issues with Grace-Hopper Relevant to Auto Offload
In this part of the paper, we discuss a few performance issues observed in the above application test, these issues reflect the immaturity of the software or hardware design in the current Grace-Hopper system, and could be fixed in the future to further improve performance and usability of automatic offload.
@@ 4.4.1. Counter-based page migration
The Hopper GPU has an access counter monitoring remote memory access, and will migrate memory page from CPU resident memory to GPU resident memory. Due to lack of access counter on the Grace CPU, data on GPU will not be migrated back to CPU. The details of the migration criteria is unknown.
Since the data migrated to GPU will not be migrated back, automatic offload with counter-based migration should work very similar to manually implemented Device First-Use policy, but we’ve already seen the slow performance of the counter-based migration in application tests. Here I present a few simple dgemm test cases with different matrix sizes, and provide a deeper understanding of the issues with counter-based migration. In the following test, matrices A(M,K), B(K,N) and C(M,N) are allocated by malloc and initialized on CPU resident memory, then multiplication $C=A\times B$ is performed at least 5 times by passing these matrices to cublasDgemm, and the access-counter should allow the matrices to be migrated to GPU resident memory. The NUMA locations of the matrices are reported after each cublasDgemm call, and runtime for each call is reported. From the NUMA location, we can infer if the data is on the CPU resident memory (NUMA 0) or GPU resident memory (NUMA 1). Four sets of inputs with different matrix sizes and shapes are tested, all use FP64 precision.
@@
When the matrix dimensions becomes M=N=K=5000, the size of all the matrices are 200.0MB, The migration becomes unstable and inconsistent from run-to-run. The tests were run many times. In most cases, only matrices A and B are migrated to GPU in the first cublasDgemm call, while C stays on the CPU no matter how many more cublasDgemm call iterations are added. Occasionally, matrices A and B stays on the CPU side in the first cycle, and only get migrated to GPU after the second call. In all runs, matrix C is never migrated to GPU.
If the square matrix dimension is further increased to M=N=K=20000, matrix size becomes 3200.0MB each, only matrix A is migrated to GPU, while B and C are always on the CPU memory no matter how many more cublasDgemm cycles are added.
The case gets strange for non-square matrix dimensions. When M=32, N=2400 and K=93536, a matrix size commonly used in my PARSEC workload, the matrix size becomes 24.0MB for A, 1795.9MB for B, and 0.6MB for C. Throughout my test, only matrix A gets migrated to GPU, while B and C are always on the CPU. This is very counter-intuitive as one would expect the bigger matrix B should at least be moved as it generates the most amount of remote memory access, but NVIDIA’s counter-based migration algorithm doesn’t agree.
From these tests, we can see that NVIDIA’s algorithm tends to migrate smaller data to GPU, this could be that the algorithm only makes the decision based on a single CUDA kernel call, i.e. moving the big data is costly comparing to that single CUDA kernel runtime, so it decides not to migrated disregard the fact that the big data gets accessed by GPU many more times later on. The current counter-based migration is unpredictable and inconsistent, NVIDIA should provide a simple way to disable it by users instead of requiring unload a kernel model by root.
4.4.2. Impact of page sizes
@@
The Grace-Hopper platform, similar to all other ARM-based platform, supports two base page sizes 4KB and 64KB. Most of the NVIDIA’s internal tests are done on 64KB page size which is also the recommended page size. At TACC, we have setup both page sizes for testings, and performance issues with the 64KB page size are revealed by comparing the test results.
Here we again run simple dgemm tests to understand the performance and issues. Notice that the aforementioned counter-based page migration mechanism doesn’t work with 4KB page size, so I can measure the performance of GPU kernel running on LPDDR5X, while I can’t do this when page size is 64KB as matrices are partially migrated to GPU as explained in Section 4.4.1.
Test results are summarized in Table 6. It can be seen that CPU accessing HBM3 memory is substantially slower under 64KB page than 4KB page for both problem sizes. There is a substantial difference even for CPU accessing LPDDR5 for the second workload, runtime with 64KB page is 15.8ms, much slower than the 10.9ms under 4KB page.
4.4.3. GPU kernel accessing system allocated HBM
As seen above in application tests, the BLAS runtime under Device First-Use policy is noticeably slower than the BLAS running on the cudaMalloc memory created in the Mem-Copy policy. Further investigation shows CUDA kernel is slow when running on system malloc allocated HBM, unless the matrices are aligned to the page. Table 7 shows the difference with page size being 64 KB. All matrices are allocated by malloc, and pinned to HBM by using numactl -m 1 ./exe. When the matrices are aligned to page size, cublasDgemm speed can be up to nearly 50% faster than if data isn’t aligned. The impact is more significant memory bandwidth bound kernels. The aligned performance is identical to the same CUDA kernel executing on cudaMalloc memory.
@@ 4.4.3. GPU kernel accessing system allocated HBM
5. Conclusion
In this paper, we report a new tool that can intercept Level 3 BLAS symbols in a CPU code, and automatically perform GPU offload using GPU-enabled BLAS along with several choices of data management strategies, especially a newly proposed first-touch type of matrix migration scheme. The tool is fine-tuned to achieve minimum overhead and take full advantage of the new unified memory architecture with cache coherent NVLink C2C. Performance tests on BLAS heavy codes show significant speedup comparing to CPU code, and the tool far outperforms NVBLAS auto-offload tool provided by NVIDIA on not only Grace-Hopper but also on the conventional PCIe-based GPU. The tool is useful for users and code developers to quickly explore the potential benefits of using GPU and have a quick start on the new architecture.
In this paper, a previous proof-of-concept BLAS auto-offload prototype tool is further optimized and extended to all level-3 BLAS operations. The limitations and advantages of the memory subsystem in NVIDIA Grace-Hopper architecture are outlined. To overcome the data transfer bottleneck between CPU and GPU, an OpenMP first-touch type of data management strategy (GPU first-use policy) is introduced and discussed in detail, the new policy minimizes the data transfer between CPU and GPU in practical scientific computing applications resulting in production level GPU performance for several quantum chemistry codes. Such data management policy can be universally applied for any GPU architecture allowing cache-coherent access to the GPU memory from CPU, eliminating the necessity of integrating a single small piece of HBM in an APU.

### v2

@@ 4.3. Application Test 2: PARSEC
PARSEC (Pseudopotential Algorithm for Real-Space Electronic Calculations)(Kronik et al., 2006; Chelikowsky et al., 1994) is a package designed to perform Density Functional Theory (DFT) calculations of solids and molecules. It solves the Kohn–Sham equations directly in real space, avoiding the use of explicit basis sets. Our benchmark case calculates energy of a Silicon nanocrystal $Si_{1947}H_{604}$, boundary sphere radius is set to 50 bohr, grid spacing is 0.9 bohr, the calculation is limited to two self-consistent field steps to reduce the benchmark cost, but the performance characteristics of a fully converged calculation are identical.
@@
In these tests, the Mem-Copy data policy resulted in runtimes slower than dual-CPU execution. The cudaMemcpy operations consumed 220 seconds, accounting for more than 50% of the total runtime — a significantly higher proportion than the one observed in the MuST test case. This is due to the fact that most matrices used in PARSEC are long skinny matrices rather than squared shape ones. For example, a common dgemm input in PARSEC is transA=’T’, transB=’N’, M=32, N=2400, K=93536, this extreme skinny shapes make the total byte size of the matrices much bigger than if square matrices are used in a calculation with equivalent computational workload. This outcome again demonstrates that the conventional data movement strategy for automatic offload is impractical even on Grace-Hoper where NVLink-C2C transfer rate is 450 GB/s per direction, not to mention the PCIe based cards where PCIe Gen5 x16 can only do 64 GB/s. The counter-based data migration strategy performs even worse due to incompetent migration algorithm, see complete discussions of NVIDIA’s migration issues later in Section 4.4.1.
Finally, the Device First-Use policy is able to efficiently manage the data movement, enabling significant performance gains. The total runtime is nearly 2x faster than the CPU run, with the dgemm component achieving nearly 10x speedup compared to the CPU run. The data transfer overhead is minimal, totaling just 1.3 seconds. The speedup is able to offset the extra cost from power hungry GPU. Again, data reuse is counted and for every matrix that is migrated to the GPU resident memory, it is reused on average 570 times by subsequent dgemm calls.
@@ 4.4. Performance Issues with Grace-Hopper Relevant to Auto Offload
In this part of the paper, I discuss a few performance issues observed in the above application test, these issues reflect the immaturity of the software or hardware design in the current Grace-Hopper system, and could be fixed in the future to further improve performance and usability of automatic offload.
@@ 4.4.1. Counter-based page migration
The Hopper GPU has an access counter that monitors remote memory access, and migrates memory pages from CPU resident memory to GPU resident memory. Due to the lack of access counter on the Grace CPU, data on GPU will not be migrated back to CPU. The details of the migration criteria are unknown.
Since the data migrated to GPU will not be migrated back, automatic offload with counter-based migration should work very similar to manually implemented Device First-Use policy, but we have already seen the slow performance of the counter-based migration in application tests. Here I present a few simple dgemm test cases with different matrix sizes, and provide a deeper understanding of the issues with the counter-based migration. In the following test, matrices A(M,K), B(K,N) and C(M,N) are allocated by malloc and initialized on CPU resident memory, then multiplication $C=A\times B$ is performed at least 5 times by passing these matrices to cublasDgemm, so that the access-counter should allow the matrices to be migrated to GPU resident memory. The NUMA locations of the matrices are reported after each cublasDgemm call, and runtime for each call is reported. From the NUMA location, we can infer whether the data is on the CPU resident memory (NUMA 0) or GPU resident memory (NUMA 1). Four sets of inputs with different matrix sizes and shapes are tested, all of which use FP64 precision.
@@
When the matrix dimensions become M=N=K=5000, the sizes of all the matrices are 200.0MB, and the migration becomes unstable and inconsistent from run-to-run. The tests were run many times. In most cases, only matrices A and B are migrated to GPU in the first cublasDgemm call, while C stays on the CPU no matter how many more cublasDgemm call iterations are added. Occasionally, matrices A and B stays on the CPU side in the first cycle, and only get migrated to GPU after the second call. In all runs, matrix C is never migrated to GPU.
If the square matrix dimension is increased to M=N=K=20000, matrix size becomes 3200.0MB each, only matrix A is migrated to GPU, while B and C are always on the CPU memory no matter how many more cublasDgemm cycles are added.
The case gets strange for non-square matrix dimensions. When M=32, N=2400 and K=93536, a matrix size commonly used in the PARSEC workload, the matrix size becomes 24.0MB for A, 1795.9MB for B, and 0.6MB for C. Throughout my test, only matrix A gets migrated to GPU, while B and C are always on the CPU. This is very counter-intuitive as one would expect the bigger matrix B should at least be moved, as it generates the largest volume of remote memory accesses, but NVIDIA’s counter-based migration algorithm fails to move it.
From these tests, we can see that NVIDIA’s algorithm tends to migrate smaller data to GPU. The reason could be that the algorithm only makes the decision based on a single CUDA kernel call, i.e. moving the big data is costly comparing to that single CUDA kernel runtime, so it decides not to migrate disregarding the fact that the big data gets accessed by the GPU many more times later on. The current counter-based migration is thus unpredictable and inconsistent, and NVIDIA should provide a simple way to disable it by users instead of requiring to unload a kernel model by root.
4.4.2. Impact of page sizes on counter-based migration
@@
The Grace-Hopper platform, similar to all other ARM-based platforms, supports two base page sizes 4KB and 64KB. Most of the NVIDIA’s internal tests are done on 64KB page size which is also the recommended page size. At TACC, both page sizes are setup for testings, and performance issues with the 64KB page size are revealed by comparing the test results.
Here I again run simple dgemm tests to understand the performance and issues. Notice that the aforementioned counter-based page migration mechanism does not work with the 4KB page size, so one can measure the performance of GPU kernel running on LPDDR5X, while this cannot be done when page size is 64KB as matrices are partially migrated to GPU as explained in Section 4.4.1.
The test results are summarized in Table 6. It can be seen that CPU accessing HBM3 memory is substantially slower under 64KB page than 4KB page for both problem sizes. There is a substantial difference even for CPU accessing LPDDR5 for the second workload, runtime with 64KB page is 15.8ms, much slower than the 10.9ms under 4KB page.
4.4.3. Sensitivity of page alignment for GPU accessing system allocated HBM
As seen above in application tests, the BLAS runtime under Device First-Use policy is noticeably slower than the BLAS running on the cudaMalloc memory created in the Mem-Copy policy. Further investigation shows that the CUDA kernel is slow when running on system malloc allocated HBM, unless the matrices are aligned to the page. Table 7 shows the difference with page size being 64 KB. All matrices are allocated by malloc, and pinned to HBM by using numactl -m 1 ./exe. When the matrices are aligned to page size, cublasDgemm speed can be up to nearly 50% faster than if data is not aligned. The impact is more significant in memory bandwidth bound kernels. The aligned performance is identical to that of the same CUDA kernel executing on cudaMalloc memory.
@@ 4.4.3. Sensitivity of page alignment for GPU accessing system allocated HBM
5. Conclusion and Future Work
In this paper, a previous proof-of-concept BLAS auto-offload prototype tool is further optimized and extended to all level-3 BLAS operations. To overcome the data transfer bottleneck between CPU and GPU, an OpenMP first-touch type of data management strategy, namely GPU first-use policy, is proposed and discussed in details. The new policy minimizes the data transfer between CPU and GPU in practical BLAS-heavy scientific computing codes, such as the quantum chemistry and quantum physics applications. Performance testings for several such codes show production level GPU performance. As all the performance evaluations were carried out on the NVIDIA Grace-Hopper architecture, limitations and advantages of the memory subsystem in Grace-Hopper are outlined. The proposed GPU first-use data management policy is applicable not limited to the architecture tested here, but can be universally applied for any CPU-GPU architecture that allows cache-coherent access of the GPU memory from CPU. This tool enables domain scientists to rapidly adopt modern GPUs in complex or legacy, BLAS-heavy codes to speedup scientific discoveries, and can also be used to quickly assess the potential benefits of GPU acceleration before a formal effort of porting. Work is in progress is to support auto BLAS offload on AMD GPUs. Additionally, FFTW auto-offload is under exploration.